# SE(3)-Equivariant and Noise-Invariant 3D Motion Tracking in Medical Images

### v1

@@ 5.2 Sensitivity and ablation studies
Table 2: Ablation study with mean (standard deviations) statistics.
Method Rot. err. (${}^{\circ}$) Trans. err. (vox.) Dice
E-CNN only 22.4 (18.3) 1.4 (1.3) .84 (.15)
@@
EquiTrack, geodesic loss 4.2 (3.8) 0.4 (0.4) .93 (.06)
@@
EquiTrack, closed-form from [35] 4.5 (6.1) 0.5 (0.8) .92 (.11)
In this section, we assess different aspects of our method, starting with its sensitivity to motion and noise. Specifically, we simulate testing pairs on Fetal-I by first using small augmentations (fixed rotation angles of ${}^{\circ}$, shifts of 2 vox., $\sigma_{B}=\sigma_{\xi}=0$) and increasing each value separately. Here we compare EquiTrack to the best variant of each baseline. Fig.4 shows that, although the learning-based methods and ANTs yield a constant accuracy across the ranges of intensity augmentations, they quickly decline with larger motion, and especially rotations for which they lose at best 13 Dice points between ${}^{\circ}$ and ${}^{\circ}$ rotations. Crucially, EquiTrack overcomes this issue and maintains high accuracy across all augmentations, with all Dice scores above 0.94. This considerable improvement further highlights the benefit of using an SE(3)-equivariant network, which enables the identification of robust spatial features across a large range of poses.
We then conduct an architectural ablation study to evaluate the design of our method. Here, we start from the E-CNN $\Phi$ (i.e., the model used in our first work [14]) and reconstruct EquiTrack by successively adding data augmentation (i.e., $\zeta$) for the inputs of $\Phi$, and the denoiser $\Psi$. The results, computed on the fetal pairs of 5.1 and reported in Table 2, first expose that steerable E-CNNs alone cannot correctly register noisy fetal data (${}^{\circ}$ angle error). Further, steerable E-CNNs have a limited learning capacity, since augmenting the inputs of $\Phi$ with the noise operator $\zeta$ only yields minor improvements and remains highly inaccurate (${}^{\circ}$ angle error). In contrast, adding the denoiser $\Psi$ (trained to undo $\zeta$ augmentations) considerably enhances performance, as it provides the E-CNN with noise-free inputs that approximately satisfy $I^{m}=T\circ I^{f}$. Lastly, we re-assemble all components, where using $\zeta$ to also train $\Phi$ leads to small but computationally cheap improvements (${}^{\circ}$, 0.1 vox.).
We also observe that the supervised geodesic loss of [7] is less robust than the L2 image loss (standard deviation higher by 3 Dice points). Further, ablating the channel weights in our closed-form algorithm highlights the benefit of this mechanism (3 Dice points and ${}^{\circ}$ in angle error). Lastly, replacing our closed-form algorithm with the KeyMorph variant yields more unstable results (+6 Dice points in standard deviations) as the latter does not account for potential landmark misalignments.
5.3 Testing on real fetal time series
Table 3: Mean (std.) statistics on real time series from Fetal-I and Fetal-II. Km-svd is short for KeyMorph-svd. Best scores are in bold and marked with * if significant at 5% (Bonferroni- corrected non-parametric Wilcoxon signed-rank test).
Method Rot. error (${}^{\circ}$) Trans. error (vox.) Dice
@@ 5.3 Testing on real fetal time series
Figure 6: Illustration of brain motion tracking for a clinical time series from Fetal-II. We show the results of ANTs, KeyMorph-SVD (KM) and EquiTrack. All time frames (top) are rigidly registered to the very first one (Reference). We overlay the reference brain mask on the results for visual assessment. Major mistakes are marked with arrows. While motion tracking was performed on masked scans, we propagate the estimated transforms to the unmasked scans to better visualise the scale of involved movements. Note how the fetus only starts substantially moving in the last four frames.
@@
Table 3 presents the results, confirming the robustness of EquiTrack, whose scores are very similar to the previous experiments despite more complex contrast changes between frames. Further, EquiTrack only loses 2 Dice points when tested on the Fetal-II dataset (unseen during training), which highlights its generalisation across sequences (contrast and resolution) and populations (healthy vs. diseased). Importantly, EquiTrack significantly outperforms all baselines on both datasets for all metrics, except for ANTs that is better by 0.2 vox. for translations on Fetal-I.
Interestingly, the performance gap between EquiTrack and its competitors is smaller than in previous experiments. To better understand this result, we plot Dice scores as a function of motion (Fig.5), which reveals two findings. First, fetuses remain stationary most of the time, i.e., with motion below ${}^{\circ}$/2 vox. for rotations/translations. In this case, all methods are fairly accurate. However, we observe an increasing gap between our method and other approaches as motion magnitude increases (Fig.6), confirming the results of the sensitivity analysis. For example, for angles above ${}^{\circ}$, EquiTrack is better than KeyMorph-SVD by 8 Dice points on Fetal-I, and superior to ANTs by 7 Dice points on Fetal-II. Overall, EquiTrack outperforms all baselines, especially for larger movements, which are the main challenge in tracking fetal motion.
@@ 6 Discussion
Since regular CNNs are much better at learning noise invariance, we introduce a denoising network to map noisy inputs to a common noise-free space. Hence, denoised volumes only differ by the pose of the represented object, such that we can directly exploit the equivariance of the E-CNN to extract corresponding anatomical features. The proposed approach can be seen as a way to decouple the processing of intensity and anatomical features. These features can eventually be registered by computing summary statistics (i.e., centres of mass) that we align with a closed-form algorithm. Importantly, this algorithm is derived to minimise the overall distance between corresponding points, which makes it robust against slight misalignments that might have propagated through our framework. As a result, our method accurately tracks motion in challenging fetal MRI time series, even under substantial domain shifts in sequence, resolution, and population.
@@
While the proposed method yields state-of-the-art results in rigid motion tracking, it has limitations. First, the E-CNN expects the two images to be the same up to an unknown rigid transform. While we mitigate this constraint by using a denoiser for motion tracking, where we register the same object seen under different poses, this is a much harder problem for inter-subject registration. Hence, our method has not been validated on general rigid registration.
@@ 7 Conclusion
We have presented a novel method based on equivariant networks for motion tracking. We demonstrated EquiTrack in the challenging scenario of tracking clinical fetal time series, where it outperformed state-of-the-art learning methods and ANTs, while running an order of magnitude faster. The enhanced tracking accuracy and robustness were achieved by leveraging a two-step architecture to first remove anatomically irrelevant sources of noise in the inputs and then compute equivariant features for accurate motion prediction.
Future work will first focus on extending our approach to general rigid registration, possibly by combining standard and equivariant filters [69] to simultaneously learn noise invariance and spatial equivariance. Further, while we track the fetal head, new piecewise-rigid equivariant models may be beneficial to track the entire fetal body and uterus [70]. Another direction is to improve our representation of spatial features by diversifying the outputs of the E-CNNs with higher-order fields, which will necessitate novel robust statistics. Finally, we will deploy our method for prospective slice prescription in 2D high-resolution fetal acquisitions, where fast 3D EPI sequences can be acquired to track motion between slices.

### v2

@@ 5.2 Sensitivity and ablation studies
In this section, we assess different aspects of our method, starting with its sensitivity to motion and noise. Specifically, we simulate testing pairs on Fetal-I by first using small augmentations (fixed rotation angles of 15, shifts of 2 vox., $\sigma_{B}=\sigma_{\xi}=0$) and increasing each value separately. Here we compare EquiTrack to the best variant of each baseline (ANTs for optimisation strategies, 3D-Pose-Net, DLIR, and KeyMorph-SVD), as well as an E-CNN alone (i.e., without a denoiser). Fig.6 shows that the accuracy of learning methods quickly declines with larger motion, and especially rotations for which they lose at least 13 Dice points across the [15, 90] range. In contrast, the E-CNN alone maintains a very high accuracy across the studied ranges of translations and rotations, due to its SE(3)-equivariance. However, the E-CNN alone is not able to cope with increasing intensity augmentations (losing 24 and 13 Dice points in the bias field and noise sensitivity studies, respectively). On the other hand, the regular CNNs used in learning methods yield constant accuracy across the ranges of simulated bias fields and noises, which highlights their ability to learn noise invariance when trained with appropriate intensity augmentations, and confirms the results of the previous experiment (inefficiency of the denoiser in KeyMorph-denoiser). Crucially, EquiTrack is able to get the best of both worlds by combining the rigid equivariance of the E-CNN with the noise invariance of the denoising CNN, and yields scores above 94 Dice points in all cases.
Table 2: Ablation study on Fetal-I with mean ($\pm$ standard dev.) scores.
Method Rot. err. () Trans. err. (vox.) Dice
E-CNN alone 22.4 (18.3) 1.4 (1.3) .84 (.15)
@@
EquiTrack, closed-form from [36] 4.5 (6.1) 0.5 (0.8) .92 (.11)
@@
Figure 7: Sensitivity study on Fetal-I (mean $\pm$ standard deviation) for the number of SE(3)-equivariant layers and number $K$ of output feature maps of the pose regressor $\Phi$. Red indicates the default values, which were selected on an independent validation set (see Section 3.6.2).
We then conduct an architectural ablation study. Here, we start from the E-CNN $\Phi$ (i.e., the model used in our first work [14]) and reconstruct EquiTrack by successively adding data augmentation (i.e., $\zeta$) for the inputs of $\Phi$, and the denoiser $\Psi$. Results are computed on the simulated fetal pairs of 5.1 (i.e., with intensity variations). Table 2 first expose that steerable E-CNNs alone cannot correctly register noisy fetal data (22.4 angle error). Further, steerable E-CNNs have a limited learning capacity, since augmenting the inputs of $\Phi$ with the noise operator $\zeta$ only yields minor gains and remains highly inaccurate (20.2 angle error). In contrast, adding the denoiser $\Psi$ (trained to undo $\zeta$ augmentations) largely enhances performance, as it provides the E-CNN with noise-free inputs that approximately satisfy $I^{m}\!=\!T\circ I^{f}$. Finally, we re-assemble all components, where using $\zeta$ to also train $\Phi$ leads to small but computationally cheap improvements (0.1, 0.1 vox.).
We observe that replacing our closed-form algorithm with the KeyMorph variant yields more unstable results (+6 Dice points in standard deviation) as the latter does not account for potential landmark misalignments. Further, ablating the channel weights in our closed-form algorithm highlights the benefit of this mechanism (2.3 in angle error).
Finally, we study the performance of EquiTrack as a function of the number of equivariant layers and output feature maps of $\Phi$. Figure 7 reveals that our method remains robust to architectural changes, with mean results varying by less than 0.8 Dice points across the ranges of studied values. The gradually decreasing scores when using more than 5 SE(3)-equivariant layers indicate that the limited learning capacity of E-CNNs is not a problem of number of learnable parameters.
Table 3: Mean ($\pm$ std.) statistics on real time series from Fetal-I and Fetal-II. Km-svd is short for KeyMorph-svd. Best scores are in bold and marked with * if significant at 5% (Bonferroni- corrected non-parametric Wilcoxon signed-rank test).
Method Rot. error () Trans. error (vox.) Dice
@@
Figure 9: Illustration of brain motion tracking for a clinical time series from Fetal-II for ANTs, KeyMorph-SVD (KM) and EquiTrack. All time frames (top) are rigidly registered to the very first one (Reference). We overlay the reference brain mask on the results for visual assessment. Major mistakes are marked with arrows. While motion tracking was performed on masked scans, we show the results on the unmasked data to better visualise the scale of involved movements. Note how the fetus only starts substantially moving in the last four frames.
5.3 Testing on real fetal time series
@@
Table 3 presents the results, confirming the robustness of EquiTrack, whose scores are very similar to the previous experiments despite more complex contrast changes between frames. Further, EquiTrack only loses 2 Dice points when tested on the Fetal-II dataset (unseen during training), which highlights its generalisation across sequences (contrast and resolution) and populations (healthy vs. diseased). Importantly, EquiTrack significantly outperforms all baselines on both data- sets for all metrics (except ANTs for translations on Fetal-I).
Interestingly, the performance gap between EquiTrack and its competitors is smaller than in previous experiments. To better understand this result, we plot Dice scores as a function of motion (Fig.8), which reveals two findings. First, fetuses remain stationary most of the time, i.e., with motion below 15/2 vox. for rotations/translations. In this case, all methods are fairly accurate. However, we observe an increasing gap between our method and other approaches as motion magnitude increases (Fig.9), confirming the results of the sensitivity analysis. For example, for angles above 30, EquiTrack is better than KeyMorph-SVD by 8 Dice points on Fetal-I, and superior to ANTs by 7 Dice points on Fetal-II. Overall, EquiTrack surpasses all baselines, especially for larger movements, which are the main challenge in tracking fetal motion.
@@ 6 Discussion
Since regular CNNs are much better at learning noise-robustness when trained with appropriate data, we introduce a denoising network to map noisy inputs to a common noise-free space. Hence, denoised volumes only differ by the pose of the brain, such that we can directly exploit the equivariance of the E-CNN to extract corresponding anatomical features. Our approach can be seen as a way to decouple the processing of intensity and anatomical features. These features can eventually be registered by computing summary statistics (i.e., centres of mass) that we align with a closed-form algorithm. Importantly, this algorithm is derived to minimise the overall distance between corresponding points, which makes it robust against slight misalignments that might have propagated through our framework. As a result, our method accurately tracks motion in challenging fetal MRI time series, even under substantial domain shifts in sequence, resolution, and population.
@@
While the proposed method yields state-of-the-art results in rigid motion tracking, it has limitations. First, the E-CNN expects the two images to be the same up to an unknown rigid transform. While we mitigate this constraint by using a denoiser for motion tracking, where we register the same object seen under different poses, this is a much harder problem for inter-subject registration. Hence, our method has not been validated on general rigid registration. Further, by focusing on rigid motion, EquiTrack does not model non-linear local distortions caused by EPI sequences. However, this is not a problem in practice, since the scale of these distortions remains marginal compared to that of the rigid movements.
@@
Finally, we note the dependence of EquiTrack on the availability of a skull stripping method. However, we highlight that this dependence is not a problem specific to EquiTrack, since it is also a prerequisite for all baselines [7, 36, 34]. Moreover, skull stripping methods are readily available for adult brain MRI, and are already implemented in many fetal brain analysis pipelines [2, 71, 72, 7].
@@ 7 Conclusion
We have presented a novel method based on equivariant networks for motion tracking. We demonstrated EquiTrack in the challenging scenario of fetal clinical time series, where it outperformed state-of-the-art learning methods and ANTs, while running an order of magnitude faster. The enhanced tracking accuracy and robustness were achieved by leveraging a two-step architecture to first remove anatomically irrelevant sources of noise in the inputs and then compute equivariant features for accurate motion prediction.
Future work will first focus on extending our approach to general rigid registration, possibly by combining standard and equivariant filters [73] to simultaneously learn noise invariance and spatial equivariance. Further, while we focus on the fetal head, new piecewise-rigid equivariant models may be beneficial to track the whole body [74]. More generally, future work will seek to tackle affine registration by incorporating scaling equivariance [75]. Preliminary methods have been proposed for scale-roto-translation equivariance [76], but making them tractable for large 3D volumes remains challenging. Another direction is to improve our representation of spatial features by diversifying the outputs of the E-CNNs with higher-order fields, which will need novel robust statistics. We will also seek to improve our augmentation model of EPI intensity variations by directly modelling them in closed form or by learning in a data-driven way. Finally, we will address the technical practicalities of deploying EquiTrack for slice prescription in 2D high-resolution fetal scans, where fast EPI volumes can be acquired to track inter-slices motion, possibly with automated detection of low quality scans [32, 77].
@@
[…]
8 Evaluation of fetal skull stripping
Skull stripping (i.e., brain segmentation) is a crucial preprocessing step for our pipeline and the baselines [7, 34, 36], because it enables the removal of the non-rigid organs surrounding the brain. Although automated skull stripping is commonly performed for the analysis of fetal brain MRI [2, 71, 72, 7, 78], existing methods have been less extensively evaluated than for the adult brain. Therefore, this section assesses the automated fetal brain segmentation method used in this paper for preprocessing.
Here, fast automated fetal skull stripping is performed by training a segmentation UNet [58]. This network comprises 4 levels, each with 2 convolution layers of $3\!\times\!3\!\times\!$ 3 × 3 kernels, ReLU activations, and batch normalisation [59]. The first layer includes 16 feature maps, this number is doubled after each max-pooling, and halved after each upsampling. Predictions are obtained from a final convolution layer with a sigmoid activation. This network is trained with the same intensity and spatial augmentation as the denoiser $\Psi$, to which we add non-linear augmentation [53] and random flipping in all directions.
This segmentation network is trained, validated, and tested on a subset of images from Fetal-I for which we have manual ground truth brain segmentations. Importantly, we keep the same subject splits as in Section 4.1, which provides us with 74 (training), 13 (validation), and 39 (testing) images from 21, 5, and 12 time series, respectively. Moreover, we also include 73 testing images with ground truth from Fetal-II to evaluate the performance of the segmenter under domain shift.
Table A1: Mean (std.) and worst-case Dice scores obtained by our skull stripping network for the preprocessing step.
Fetal-I Fetal-II
No dilation 4-voxel dilation No dilation 4-voxel dilation
Mean Dice 0.93 (0.02) 0.94 (0.01) 0.91 (0.02) 0.93 (0.02)
Worst-case Dice 0.88 0.91 0.86 0.89
Table A1 shows that the segmenter produces high quality segmentations on Fetal-I with average and worst-case scores of 93 and 88 Dice points, respectively. Moreover, we see that our 4-voxel dilation strategy (which is applied to both the predicted and ground truth segmentations) further improves results (worst-case 91 Dice points) by compensating for small under-segmented areas in the original predictions. Finally, the segmenter is also robust against slight domain gaps, since it yields accurate scores on Fetal-II (worst-case 89 Dice points) when faced with intensity and morphological shifts.